OpenStudy (anonymous):

Please help me solve this! Only looking for help, not answers! Equation posted below:

OpenStudy (anonymous):

\[\frac{ 3x+6 }{ 8} = \frac{ 7x-1 }{ 6 }\]

OpenStudy (anonymous):

cross multiply and simplify

OpenStudy (anonymous):

\[\frac{ 24 }{ 1 } (3x+6) = \frac{ 42 }{ 1}(7x-1) ???\]

OpenStudy (anonymous):

6(3 x+6)=8(7x-1)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

18x+36= ?

OpenStudy (anonymous):

18x=36 = 56x-2??

OpenStudy (anonymous):

Oops, sorry, made a typo. 18x+36

OpenStudy (anonymous):

18x+36=56x-8 transpose 56x to left hand side and 36 to right hand side. then find x

OpenStudy (anonymous):

18 x-56 x=-8-36 -38 x=-44 x=?

OpenStudy (anonymous):

x=1.15?

OpenStudy (anonymous):

\[x=\frac{ -44 }{ -38 }=\frac{ 22 }{ 19 }\]

OpenStudy (anonymous):

Oh... I'm so confused. Math and I just don't get along well, haha! Thank you for the help by the way! :)

OpenStudy (anonymous):

yw